I need help with this inequality.
3(5-x)≥7x-1

Answer by sofiyac(983)

15-3x>=7x-1 add 3x to each side
15>=10x-1 add 1 to each side
16>=10x divide each side by 10
16/10>=x reduce top and bottom by 2
8/5>=x